### v1

@@ 4.3 Discussion on Patch-level Descriptors
Run-time & Query Speed. As we reduce the dimensionality of the DINO v2 features to 256, the embeddings of all patches of each image need a total of 512 kB, while, for AnyLoc, the storage per image is 384 kB. While we require slightly more storage than AnyLoc [33], the difference is small. In addition, we compare the time of querying an image from the database of different sizes using AnyLoc or VOP in seconds. Prefiltering top 20 images by DINO V2 CLS tokens and running VOP for reranking to get top-1 out of 500 images cost 0.009 seconds, while 0.003 for AnyLoc. Querying top-1 from 5k images (prefiltered to 100), our method costs 0.03 seconds, while AnyLoc runs in 0.02s. Note that VOP needs much more time for radius search without prefiltering. We recommend using VOP as a reranking method in retrieval.
@@ 5 Conclusion
Our paper introduces a novel Visual Place Recognition (VPR) method focusing on patch-level overlap prediction to address challenging view changes and occlusions. Through extensive testing on the MegaDepth, PhotoTourism and ETH3D datasets, we demonstrate the effectiveness of the proposed VOP method. It outperforms state-of-the-art retrieval approaches in terms of relative pose accuracy. Critically, we demonstrate the inadequacy of recall as the sole metric for VPR performance, advocating for metrics aligned with real-world applications. By comparing configurations optimized for AUC score and recall, our results underline the necessity of metric selection tailored to specific VPR tasks. We believe this research marks a significant step forward in VPR, emphasizing the importance of patch-level analysis and metric consideration for improved accuracy and application relevance.

### v2

@@ 4.3 Discussions
Storage & Query Speed. As we reduce the dimensionality of the DINOv2 features to 256, the embeddings of all patches of each image need a total of 512 kB, while, for AnyLoc, the storage per image is 384 kB. While we require slightly more storage than AnyLoc [25], the difference is small. Compared to storing the local features in the reranking-based methods, VOP costs less. In addition, we compare the time of querying an image from the database of different sizes using AnyLoc or VOP in seconds. Prefiltering top 20 images by DINOv2 [CLS] token and running VOP for reranking to get top-1 out of 500 images cost 0.009 seconds, while 0.003 for AnyLoc. Querying top-1 from 5K images (prefiltered to 100), our method costs 0.03 seconds, while AnyLoc runs in 0.02s. Note that VOP needs much more time for radius search without prefiltering. We recommend using VOP as a reranking method in retrieval.
Image Patching. We investigated the number of patches to be used to split the images. The experiments were conducted on the testing scenes of MegaDepth, with all the images resized to $224\times 224$. From the trained VOP model with a patch size of $14^{2}$, we can extract 256 patch descriptors. Then, average pooling is applied to aggregate the patch descriptors to different patch sizes, such as $28^{2}$, $56^{2}$, $112^{2}$, and $224^{2}$. For example, patch size = 224 will lead to a single patch of an image. The retrieval is done on the same prefiltered image list as discussed. As shown in Table 6, the aggregated patches e.g., patch size=$224^{2}$, perform worse than $14^{2}$ on MegaDepth pose estimation and could not generalize well on Inloc localization. This demonstrates that patch-level features can potentially improve estimated pose and other geometric problems.
Patch size AUC@10 $\uparrow$ med. pose err. $\downarrow$ recall@5, 0.5m    $\uparrow$
MegaDepth - top5 DUC1 DUC2
$224^{2}$ 67.0 2.09 30.8 24.4
$112^{2}$ 67.0 2.09 38.4 38.9
$56^{2}$ 66.5 2.17 48.5 57.3
$28^{2}$ 65.9 2.29 59.1 72.5
$14^{2}$ 67.6 2.03 72.2 77.1
Table 6: Ablations on different patch sizes used in inference time. We show the AUC@10 scores and median pose errors of the top-5 retrieved images on MegaDepth [49], and the recall@, 0.5m on the localization data, Inloc [51] (DUC1, DUC2).
Qualitative Results. Most VPR methods prioritize retrieving similar images, typically resulting in short baselines that are not suitable for reconstruction. These goals conflict: the most similar images often produce short baselines, making pose estimation unstable. We aim to move beyond traditional similarity metrics and design retrieval methods tailored for geometric challenges, such as selecting images suitable for pose estimation. We visualize three query examples in Fig. 8 with their top-1 retrieved images using different methods. VOP results in low pose errors as we find images with reasonable baselines for stable pose estimation.
5 Ablations
We investigate the influence of image patches, different patch sizes, the threshold used in radius search and the influence of the data augmentation on test accuracy and generalization capability.
Radius Search Threshold. One of the parameters of the proposed VOP is the radius search threshold when selecting similar patches in the embedding space. To understand how sensitive VOP is to the setting of this threshold, we show tuning results on the validation set of the MegaDepth dataset. We show the results with different thresholds (horizontal axis) on the validation set in Fig. 7. While the accuracy varies with the threshold, setting it on the validation set and using this fixed threshold on all tested large-scale datasets leads to SOTA performance. High AUC scores and low pose errors are preferred, thus, threshold = 0.15 is chosen, i.e., the patches with the descriptor similarities lower than 0.15 are recognized as neighbors.
Dropout Layer. Table 7 shows the relative pose estimation performance on the top-$k$ images retrieved by the model with the searching radius and votings chosen from the best AUC, with or w/o global prefilter ([CLS] tokens), and dropout layer. Testing sets of MegaDepth are used. The CLS prefiltering improves the AUC scores. However, it increases the median pose errors marginally at the same time. As shown in the fifth row, data augmentation is essential in robustly learning the embeddings. Also, the last two rows in Table 7 show that the dropout layer improves the performance on MegaDepth and helps with generalization for pose estimation on other data and indoor localization.
Prefilter Augment Dropout AUC@10 $\uparrow$ Med. pose error ($\downarrow$ inliers $\uparrow$
$\checkmark$ $\times$ $\times$ 65.1 2.19 272.5
$\times$ $\checkmark$ $\times$ 66.3 2.18 222.0
$\checkmark$ $\checkmark$ $\times$ 66.7 2.18 263.0
$\checkmark$ $\checkmark$ $\checkmark$ 67.6 2.02 246.5
Table 7: Relative pose estimation on the MegaDepth dataset [49] on the top 5 retrieved images using different configurations, with the best results in bold. Prefilter indicates if the [CLS] token was employed to shortlist the potential candidates before overlap prediction. Augment refers to whether data augmentation was used.
Figure 7: Ablations on the threshold used in radius search. AUC@10 and median pose errors on the validation scene shown.
Figure 8: Baselines and pose errors between the retrieved images and queries using different methods shown.
@@ 6 Conclusion
We introduce a novel Visual Place Recognition method, VOP, focusing on patch-level visual overlap prediction to address challenging viewpoint changes and occlusions. The proposed VOP predicts overlapping image patches by learning a patch-level descriptor and employing a robust voting mechanism and radius search in the descriptor space. Through extensive testing on the MegaDepth, PhotoTourism, ETH3D, and InLoc datasets, we demonstrate the effectiveness of the proposed VOP. On average, it achieves the highest accuracy, improving upon the state-of-the-art AnyLoc by two AUC points and reducing its median pose error by 25%. We believe this research marks a significant step forward in VPR, breaking the frame of image similarity prediction (through global or local features) to obtain a patch-level understanding of the query and database images.